OpenStudy (anonymous):

Why does n^2 = (n-1)^2 + (n-1) + n

OpenStudy (anonymous):

left hand side of the equation is = (n-1)^2+(n-1)+n =n^2-2n+1+n-1+n =n^2-2n+2n+1-1 =n^2 the right hand side

OpenStudy (anonymous):

thanks